Arya’s Alchemy produces mineral based face masks. The company has the following financial information:

The company’s sales price is $20 per unit. The variable costs of producing masks is $6 per unit. The company expects to have fixed costs of $10,000 next year. The company expects to sell 1,000 masks. Assume no taxes.

Calculate the breakeven point in units: ____ units need to be sold.

Calculate the breakeven point in dollars: ___ worth of product need to be sold.

How many units must the company sell to reach a target profit of $25,000?

Prepare a budgeted contribution format income statement. The contribution margin is: ___ and the operating income is: ___

Compute the margin of safety in both dollar and percentage terms. ____ is

the dollar amount sales could drop before the company starts to lose money. __  is the margin of safety as

Compute the degree of operating leverage: _____

  • If sales increase by 20% in the following year, how much would net income increase (use the degree of operating leverage to compute your answer). Net income would increase by ____ and the new net income would be _____
Answer #1

a) Break even unit = 10000/(20-6) = 714 Units

Break even sales = 714*20 = $14280

b) required unit = (10000+25000)/14 = 2500 Units

Contribution margin income statement

Sales 20000
Variable cost 6000
Contribution margin 14000
Fixed cost 10000
Operating income 4000

Margin of safety = 20000-14280 = 5720

Margin of safety (%) = 5720/20000 = 28.6%

Degree of operating leverage = 14000/4000 = 3.5

Net income would increase by (20*3.5) = 70%

New net income = 4000*1.7 = 6800